Section 04  Algebraic Geometry

Section 04: Algebraic Geometry

Invited Lectures

Wednesday, 19 August  MA 005

16:15 – 17:00 Robbert Dijkgraaf (also Section 11), University of Amsterdam, Amsterdam, The Netherlands The Mathematics of Fivebranes (I: p.62) Chair: Fedor Smirnov

17:15 – 18:00 Johan de Jong, Princeton University, Princeton, USA Barsotti-Tate groups and crystals (I: p.61) Chair: Gerard van der Geer

Thursday, 20 August  MA 005

14:00 – 14:45 Mark L. Green, UCLA, Los Angeles, USA Higher Abel-Jacobi Maps (I: p.62) Chair: Frank Schreyer

15:00 – 15:45 Maurizio Cornalba, Università di Pavia, Pavia, Italy Cohomology of moduli spaces of stable curves (I: p.61) Chair: Alberto Conte

Friday, 21 August  MA 005

14:00 – 14:45 Mikhail M. Kapranov, Northwestern University, Evanston, USA Operads and algebraic geometry (I: p.65) Chair: Enrico Arbarello

15:00 – 15:45 Victor Vadimovich Batyrev, Universität Tübingen, Tübingen, Germany Mirror symmetry and toric geometry (I: p.60) Chair: Yujiro Kawamata

Saturday, 22 August  H 1028

16:15 – 17:00 Paul S. Aspinwall (also Section 11), Duke University, Durham, USA Geometry and String Duality (I: p.59) Chair: Pelham M.H. Wilson

Algebraic Geometry  Section 04

Short Communications and Poster Sessions

Short Communications  Thursday, 20 August
Toric Varieties  MA 042

Chair: Klaus Altmann

16:15 – 16:30 Anette A’Campo-Neuen, Universität Konstanz, Konstanz, Germany Quasi–Projective Reduction of Toric Varieties (p.49)

16:30 – 16:45 Jürgen Hausen, Universität Konstanz, Konstanz, Germany Toric Quotients (p.53)

16:45 – 17:00 Evgeni Nikolaevich Materov, Krasnoyarsk State University, Krasnoyarsk, Russia One estimate for dimension of H^{n}(C^{n}\V, C) (p.56)

Singularities I  MA 042

Chair: Gert-Martin W. Greuel

17:15 – 17:30 Sabir M. Gusein-Zade, Moscow State University, Moscow, Russia On the index of a gradient vector field at an isolated complete intersection singularity (p.53)

17:30 – 17:45 Ludwig Balke, Universität Bonn, Bonn, Germany Open book fibrations of lens spaces and deformations of cyclic quotient singularities (p.50)

17:45 – 18:00 Claus Hertling, Universität Bonn, Bonn, Germany Variational Torelli result for hypersurface singularities (p.53)

Short Communications  Friday, 21 August
D-Modules/Vector Bundles I  MA 042

Chair: Sabir M. Gusein-Zade

16:15 – 16:30 Aleksandr G. Aleksandrov, Institute of Control Sciences, Moscow, Russian Federation Gauß-Manin-Zusammenhang und logarithmische Picard-Fuchs-Systeme (p.49)

16:30 – 16:45 Vladimir Bavula, University of Edinburgh, Edinburgh, UK The Simple Holonomic Modules over the Ring of Differential Operators on a Smooth Algebraic Surface (p.50)

16:45 – 17:00 Ushadevi N. Bhosle, Tata Institute of Fundamental Research, Mumbat, India Moduli spaces of torsion-free sheaves on nodal curves (p.51)

Deformations and Degenerations  MA 042

Chair: Aleksandr G. Aleksandrov

17:15 – 17:30 Kimio Miyajima, Dept. of Math. and Comp. Science, Kagoshima City, Japan CR construction of the semi-universal family of normal isolated singularities (p.57)

17:30 – 17:45 Eduardo S. Esteves, IMPA, Rio de Janeiro, Brazil Degenerations of Weierstrass points (p.52)

17:45 – 18:00 Diego Paccagnan, Universitá di Padova, Padova, Italy Maximum number of cusps on algebraic plane curves and nodes on surfaces (p.59)

Short Communications  Saturday, 22 August
Affine Algebraic Geometry  MA 042

Chair: Marko Roczen

14:00 – 14:15 Tatsuji Kambayashi, Tokyo Denki University, Saitama-Ken, Japan A Nullstellensatz for Pro-Affine Algebras (p.55)

14:15 – 14:30 Vladimir Shpilrain, The University of Hong Kong, Hong Kong, P.R. of China Embeddings of curves in the plane (p.60)

14:30 – 14:45 Jie-Tai Yu, The University of Hong Kong, Hong Kong, P.R. of China Polynomial Retracts and the Jacobian Conjecture (p.63)

Section 04  Algebraic Geometry

Riemann Surfaces/Linear Systems  MA 042

Chair: Anette A’Campo-Neuen

15:00 – 15:15 Sergei Natanzon, Independent University of Moscow, Moscow, Russia Topological structure of spaces of holomorphic morphisms (p.58)

15:15 – 15:30 Masahiro Ohno, Waseda University, Dept. of Math., Tokyo, Japan On nef values of determinants of ample vector bundles (p.59)

15:30 – 15:45 Siye Wu, The University of Adelaide, Adelaide, SA, Australia On the instanton complex of holomorphic Morse theory (p.62)

Curves and Surfaces/Higher-Dimensional Varieties I  MA 042

Chair: Rolf-Peter Holzapfel

17:15 – 17:30 Viktor S. Kulikov, Steklov Mathematical Institute, Moscow, Russia On Chisini’s Conjecture (p.55)

17:30 – 17:45 Kazuhisa Maehara, Tokyo Institute of Polytechnics, Tokyo, Japan Deformation of function fields (p.56)

17:45 – 18:00 Dan Avritzer, Universidade Federal de Minas Gerais, Belo Horizonte, Brazil The Hilbert scheme component of the intersection of two quadrics (p.50)

Short Communications  Monday, 24 August
Moduli Spaces and Mathematical Physics  MA 042

Chair: Ragni Piene

14:00 – 14:15 Sheldon H. Katz, Oklahoma State University, Stillwater, OK, USA Geometric Engineering of Quantum Field Theories (p.55)

14:15 – 14:30 Alexandre Kabanov, Michigan State University, East Lansing, MI, USA Intersection theory on the moduli spaces of stable curves (p.54)

14:30 – 14:45 Tyler J. Jarvis, Brigham Young University, Provo-UT, USA Geometry of the Moduli of Higher Spin Curves (p.54)

Vector Bundles II  MA 042

Chair: Herbert Kurke

15:00 – 15:15 Hagen Meltzer, Technische Universität Chemnitz, Chemnitz, Germany Exceptional vector bundles (p.57)

15:15 – 15:30 Vasile P. Brinzǎnescu, Institute of Mathematics, Bucharest, Romania Moduli spaces of stable 2-vector bundles over ruled surfaces (p.51)

15:30 – 15:45 Elham Izadi, University of Georgia, Athens, Georgia, USA An embedding of a moduli space of vector bundles (p.54)

Singularities II  MA 309

Chair: Herbert Kurke

16:15 – 16:30 Michał Kwieciński, Jagiellonian University, Kraków, Poland An algebraic map with nonconstructible topology

Short Communications  Tuesday, 25 August
Surfaces/Higher-Dimensional Varieties II  MA 042

Chair: Vasile P. Brinzǎnescu

16:15 – 16:30 Alberto Conte, Università di Torino, Torino, Italy On the unirationality of quartic hypersurfaces (p.52)

16:30 – 16:45 Fabrizio M.E. Catanese, G.A. Universität Göttingen, Göttingen, Germany Bidouble covers and surface geography (p.51)

16:45 – 17:00 Remo Gattazzo, Università di Padova, Padova, Italy Almost factorial quartic surfaces in ℙ_{ℂ}^{3} with a tacnodal point (p.52)

Algebraic Geometry  Section 04

Surfaces/Higher-Dimensional Varieties III  MA 042

Chair: Fabrizio M.E. Catanese

17:15 – 17:30 Marina R. Marchisio, Università di Torino, Torino, Italy Some new examples of smooth unirational quartic threefolds (p.56)

17:30 – 17:45 Uwe Nagel, University-GH Paderborn, Paderborn, Germany Characterization of projective subschemes by locally free resolutions (p.57)

17:45 – 18:00 Tomasz Szemberg, Jagiellonian University, Cracow, Poland Higher order embeddings and Seshadri constants of abelian varieties (p.61)

Short Communications  Wednesday, 26 August
Surfaces/Higher-Dimensional Varieties IV  MA 042

Chair: Viktor S. Kulikov

14:00 – 14:15 Yuri G. Prokhorov, Moscow State University, Moscow, Russia Mori contractions of relative dimension 1 (p.59)

14:15 – 14:30 Maria Cristina Ronconi, Università degli Studi di Padova, Padova, Italy Algebraic threefolds with q_{1} = q_{2} = p_{g} = 0, P_{2i} = 1, and P_{2i+1} = 0 (p.60)

14:30 – 14:45 Ezio Stagnaro, Università di Padova, Padova, Italy Pluricanonical maps of two general type threefolds with p_{g} = 0 (p.60)

Algebraic Cycles  MA 042

Chair: Yuri G. Prokhorov

16:15 – 16:30 Manfred F.G. Streit, Johann Wolfgang Goethe Universität, Frankfurt, Germany Field of Definition and Galois Orbits of some Non–Congruence Subgroups (p.61)

16:30 – 16:45 Serguei G. Tankeev, Vladimir State University, Vladimir, Russia Algebraic cycles of small codimension on simple 2p- or 4p-dimensional abelian variety (p.62)

16:45 – 17:00 Johannes Huisman, Institut Mathematique de Rennes, Rennes, France Uniformization of real algebraic curves and complex analytic coordinates on Teichmüller spaces

(p.54)

Poster Session  Wednesday, 19 August 14:00 – 15:45  MA 144

Jacqueline Arancibia, Departmento de Matemática, UFPB, Joao Pessoa, Brasil

Canonical Curves in P^{3} (p.49) Isidro P. Nieto, Universidad Michoacana de SNH, Morelia, Mexico

H_{2,2}-Kummer surfaces associated to polarized abelian surfaces of type (1,3) (p.58) Niranjan Ramachandran, University of Michigan, Ann Arbor, MI, USA

A conjecture of Deligne (p.59) Rudi Weikard, University of Alabama at Birmingham, Birmingham, AL, USA

Algebro-geometric Differential Operators (p.62) Katsuaki Yoshida, Nihon University, Funabashi, Japan

Klein’s surface of genus three and associated theta constants (p.62) Rupert W.T. Yu, Université de Poitiers, Poitiers, France

Lattice polytopes for certain Demazure modules of sl_{n+1} (p.63)